OpenStudy (anonymous):

Using complete sentences, explain how to find the zeros of the function f(x) = 2x^3 - 9x + 3.

OpenStudy (tkhunny):

Facoring if possible. Completing the Square or the Quadratic Formula, otherwise.

OpenStudy (anonymous):

can you show me how to do it?

OpenStudy (tkhunny):

You've probably done it 100 times, already. Solve this for x: x + 2 = 0

OpenStudy (tkhunny):

Try this one: \(x^{2} - 1 = 0\)

OpenStudy (tkhunny):

How about this one: \(x^{2} - 4x + 4 = 0\)

Parth (parthkohli):

Here's how: Go to http://wolframalpha.com and type out `2x^3 - 9x + 3 = 0`. Scroll down to see the result.

OpenStudy (tkhunny):

You've been studying factoring and the Quadratic Formula. Now is the time to use these skills and tools!

OpenStudy (anonymous):

(2.*(x+2.271633299))*(x-.3422414369)*(x-1.929391862)

OpenStudy (anonymous):

so you factor to get (2.*(x+2.271633299))*(x-.3422414369)*(x-1.929391862)=0 get rid of 2 by dividing (x+2.271633299)*(x-.3422414369)*(x-1.929391862)=0 Then you just solve individually (x+2.271633299)=0 -> x=-2.27... (x-.3422414369)=0 -> x= .34... (x-1.929391862)=0 -> x=1.92...

OpenStudy (tkhunny):

Is it really \(x^{3}\) and not \(x^{2}\)?

OpenStudy (anonymous):

i think it may be x^2